George A. Madrid 28 November 1995

Critique of “The Beefeaters”

As a man who has dated a number of women with alternative eating habits (read women who wouldn’t eat beef), I got a huge kick out of this story. I really like the idea of a society that eats itself. I wonder how it came about though, and depending on the answer to this question, the story has a different effect for me.

If something has happened similar to Do Androids Dream of Electric Sheep and now animals are scarce and this is the reason no one eats beef or chickens, then I feel the scandal of what they are doing. If it is simply not socially acceptable, and even the poor will eat protein soup before eating beef, then I feel like they are crossing an artificial social line (which I find stupid being a beefeater, myself) so I feel different about them. Based on the story, I suspect the former, but I can imagine that if we stopped farming cows, they might not do too well in the wild.

I found the story a little hard to read, because there were two levels of flashbacks going on. Once or twice, I lost track of my temporal location, and was surprised later on to discover that I wasn’t where I thought I was. For example, the transition on page 6, from the past to a different point in the past with the same characters. I was expecting to be back in the present. Unless the intent is to confuse the reader a little, I think more structure needs to be provided to help the reader along.

I really enjoyed the historical background for the society of beef-eating people. This firmly locked the story to my world’s future. This made it more real for me. This had the added effect of making me wonder how the Earth went from now to then.

I appreciate that I was never actually told the difference between “meat” and “beef.” Keeping me wondering did a lot to heighten the effect.